OpenStudy (anonymous):

PLZ HELP ILL GIVE A MEDAL Christina is a 77% foul shooter. If she has to make both free throws for her team to win the game, what is the probability as a percent of her making both shots? 5,929% 0.5929% 592.9% 59.29%

OpenStudy (anonymous):

probability of winning in one throw =1-77%=23% probability of winning in both shots=23%*23%=23/100 *23/100=.0529=5.29%